If the minute hand of a big clock is 1.12m long, find the rate at which its tip is moving in centimeters per minute.
Answer by Alan3354(69443)

If the minute hand of a big clock is 1.12m long, find the rate at which its tip is moving in centimeters per minute.
---------------------
r = 112 cm
Circumference = distance/revolution = 2*pi*r/360 degs =224pi/360
1 minute (of time) = 360/60 = 6 degs
s = (224pi/360)*6
s = 56pi/15 cm
s =~ 11.729 cm
r =~ 11.729 cm/minute
----------------------
Using radians:
omega = 2pi/hour = (pi/30)rad/minute
speed = omega*r = 112pi/30 cm/min
= 56pi/15 same answer, much simpler